# DECIPHERING THE DEFINITION OF ADVERSARIAL ROBUSTNESS FOR POST-HOC OOD DETECTORS

### v3

@@ 4.1 Experiment Setup
As attack methods, we choose FGSM(-$L^{\infty}$), PGD(-$L^{\infty}$), DF(-$L^{2}$) from FoolBox [82]; and the mPGD-($L^{\infty}$). The attacked models are ResNet-18 [83], ResNet-50 [83], and Swin-T [84].
The datasets are CIFAR-10 & CIFAR-100 [22], ImageNet-1K [23] and its variant ImageNet-200 with just 200 classes.
@@ 4.3 DISCUSSION
Interestingly, several OOD benchmarks [91, 18, 5, 6] and a recent challenge [92] focus solely on natural distribution shifts, neglecting adversarial examples. This underscores a fundamental conflict between Adversarial Machine Learning (AML) and Out-of-Distribution (OOD) detection, where detectors excel either in adversarial or natural distributions, but rarely both. Current OOD detectors struggle to converge on both fronts.
Adversarial examples, as counterparts of In-Distribution (ID) data, significantly alter the attention of pre-trained classifiers, as shown in Figure 1 and discussed in Equation 3. These predominantly untargeted attacks aim to minimally change class predictions, resulting in dissimilarity in attention maps (see Grad-CAM Section 4.2). The DF attack exhibits minimal attention shifts across numerous samples, yet many attacked samples still demonstrate notable deviations in attention maps. Similarly, the mPGD attack induces perceptible perturbations, with slightly improved detection results compared to PGD, but overall performance remains insufficient.
Post-hoc detectors are trained on distributions similar to the underlying pre-trained model, yet studies (e.g., [93, 94]) indicate significant discrepancies between the training distribution and adversarial distributions. Achieving high detection rates on adversarial OOD samples is crucial for developing robust defense mechanisms, a goal that remains elusive in current research.
@@ 4. Level 4:
    Let your method fail against sophisticated attacks, such as adaptive attacks [97, 54] or design OOD adversarial examples that convert OD to ID samples [3]. Adversarial robustness is an iterative process, where defenses are proposed, evaluated, and improved upon in response to new attacks or discovered vulnerabilities.
It must be noted that not all levels are required to be evaluated. This roadmap should indicate which level of OOD detection is robust. Further research would be encouraged to do at least an adversarial robustness benchmark (level 1), even though the OOD detector is not focused on adversarial robustness.
@@ 5 CONCLUSION
In this study, we assess the performance of the 16 post-hoc OOD detectors in their ability to detect various evasive attacks. We conducted prominent white-box adversarial attacks, such as PGD and DF, on the CIFAR-10 and ImageNet-1K datasets. Our discovery indicates that current post-hoc methods are not ready for real-world applications as long as they are vulnerable against the well-known threat — adversarial examples. We hope our experiments give a baseline for further research by improving post-hoc methods towards robustness and will find a place as a standardized benchmark, such as OpenOOD.
@@
We propose to extend the experiments towards transferability because adversarial examples transfer effectively across different datasets [98] and models [99]. Then, we would suggest using black-box attacks for a realistic open-world scenario. Throughout this work, we assumed a perfect pre-trained classifier on a clean curated dataset. An imperfect pre-trained classifier would be the subject of future research [100].

### v5

@@ 4.1 Experiment Setup
As attack methods, we choose FGSM(-$L^{\infty}$), PGD(-$L^{\infty}$), DF(-$L^{2}$) from FoolBox [88]; and the mPGD-($L^{\infty}$). The attacked models are ResNet-18 [89], ResNet-50 [89], and Swin-T [90]. The datasets are CIFAR-10 & CIFAR-100 [23], ImageNet-1K [24] and its variant ImageNet-200 with just 200 classes.
@@ 4.3 DISCUSSION
Interestingly, several OOD benchmarks [97, 19, 6, 7] and a recent challenge [98] focus solely on natural distribution shifts, neglecting AdEx. This underscores a fundamental conflict between adversarial machine learning (AML) and OOD detection, where detectors excel either in adversarial or natural distributions, but rarely both. Current OOD detectors struggle to converge on both fronts.
AdEx, as counterparts of In-Distribution (ID) data, significantly alter the attention of pre-trained classifiers, as shown in Figure 1 and Equation 1. These predominantly untargeted attacks aim to minimally change class predictions, resulting in dissimilarity in attention maps (Section 4.2). The DF attack exhibits minimal attention shifts across numerous samples, yet many attacked samples still demonstrate notable deviations in attention maps. Similarly, the mPGD attack induces perceptible perturbations, with slightly improved detection results compared to PGD, but overall performance remains insufficient.
Post-hoc detectors are trained on distributions similar to the underlying pre-trained model, yet studies, e.g. [99, 100], indicate significant discrepancies between the training distribution and adversarial distributions. Achieving high detection rates on adversarial OOD samples is crucial for developing robust defense mechanisms, a goal that remains elusive in current research.
@@ 4. Level 4:
    Let your method fail against sophisticated attacks, such as adaptive attacks [102, 60] or design OOD AdEx that convert OD to ID samples [4]. Adversarial robustness is an iterative process, where defenses are proposed, evaluated, and improved upon in response to new attacks or discovered vulnerabilities.
It must be noted that not all levels are required to be evaluated. This roadmap should indicate which level of OOD detection is robust. Further research would be encouraged to do at least an adversarial robustness benchmark (level 1), even though the OOD detector has a different focus.
@@ 5 CONCLUSION
In this study, we assess the performance of the 16 post-hoc OOD detectors in their ability to detect various evasive attacks. We conducted white-box adversarial attacks, such as PGD and DF, on the CIFAR-10 and ImageNet-1K datasets. Our discovery indicates that current post-hoc methods are not ready for real-world applications as long as they are vulnerable to the well-known threat — adversarial examples (AdEx). The semantic shift was evaluated by utilizing Grad-CAM to develop a novel metric. This metric indicates that the AdEx tends to perform a minimal shift in semantics that is spatially proximate to the original, but can exhibit a wide range across the image.
We hope our experiments and datasets give a common baseline for further research by improving post-hoc methods towards robustness and will find a place as a standardized benchmark, such as OpenOOD. We hope that future post-hoc OOD detectors will include at least one level 1 evaluation, even though these detectors do not prioritize adversarial robustness.
@@
We propose to extend the experiments towards transferability because AdEx transfers effectively across different datasets [103] and models [104, 105].
Then, we would suggest using black-box attacks for a more realistic open-world scenario [106, 56]. Their attack patterns are usually easier to detect, but also transfer well to other models and datasets [107].
Throughout this work, we assumed a perfect pre-trained classifier on a clean curated dataset. An imperfect pre-trained classifier would be the subject of future research [108].